# DGP-LVM: Derivative Gaussian process latent variable models

### v2

@@ Abstract
We develop a framework for derivative Gaussian process latent variable models (DGP-LVM) that can handle multi-dimensional output data using modified derivative covariance functions. The modifications account for complexities in the underlying data generating process such as scaled derivatives, varying information across multiple output dimensions as well as interactions between outputs. Further, our framework provides uncertainty estimates for each latent variable samples using Bayesian inference. Through extensive simulations, we demonstrate that latent variable estimation accuracy can be drastically increased by including derivative information due to our proposed covariance function modifications. The developments are motivated by a concrete biological research problem involving the estimation of the unobserved cellular ordering from single-cell RNA (scRNA) sequencing data for gene expression and its corresponding derivative information known as RNA velocity. Since the RNA velocity is only an estimate of the exact derivative information, the derivative covariance functions need to account for potential scale differences. In a real-world case study, we illustrate the application of DGP-LVMs to such scRNA sequencing data. While motivated by this biological problem, our framework is generally applicable to all kinds of latent variable estimation problems involving derivative information irrespective of the field of study.
Keywords: Gaussian processes, single-cell RNA, Bayesian inference, derivatives, latent variables
@@ 1 Introduction
Gaussian processes (GPs) are a class of statistical models known for their flexible structure and favourable properties to analyze complex data (Williams and Rasmussen,, 1995). Since their inception, several extensions have been proposed, among which the most relevant ones to this paper are adding derivative information to GPs (Solak et al.,, 2002; Rasmussen and Williams,, 2006), building GPs with multiple outputs (Cressie,, 1993; Teh et al.,, 2005; Rasmussen and Williams,, 2006) as well as modeling latent input variables (Lawrence,, 2003, 2005). Since differentiation is a linear operation, any variable and its derivative would be linearly related. For the same reason, as a fundamental property of GPs, a derivative of a GP is just another GP with a related covariance function. Together, this results in a single GP model for the outputs and their derivatives with a joint derivative covariance function (Solak et al.,, 2002).
@@
In real-world data, the derivatives are seldom exactly computed, which adds a major challenge to the modeling endeavour. If derivatives are computed with respect to the observed (non-latent) inputs, they are naturally only an approximation of the derivatives with respect to latent inputs. Conversely, if the derivatives are (implicitly) computed with respect to the latent inputs, the uncertainty of the latter will induce hard-to-quantify uncertainty in the estimated derivatives. One way or another, this lack of exact derivative information poses a serious challenge for derivative GP modeling. Existing approaches are not equipped to deal with the significant scale differences between outputs and derivatives, thus requiring modifications in its covariance functions to ensure valid and efficient latent variable estimation.
In this paper, we demonstrate a combination of all the above model extensions leading to our DGP-LVM: derivative Gaussian process latent variable model framework. We provide more context about the real-world modeling challenges in Section 1.1 as a basis for our motivation to develop DGP-LVM. The remainder of the paper is structured as follows. We discuss and provide context on related works in Section 2. We introduce our methodology and model development in Section 3 and perform extensive simulation studies in Section 4 that demonstrate the relevance of our contributions. We further illustrate DGP-LVM on a real single-cell RNA sequencing data in Section 5 before discussing our methods’ limitations and future work in Section 6.
@@ 1.1 Motivation
In developmental biology, to describe temporal biological processes, researchers use stochastic approaches to understand cellular progression, that is, how cells develop and undergo changes in their state throughout various stages over the period of time (Maamar et al.,, 2007; Losick and Desplan,, 2008; Raj and Oudenaarden,, 2008). Currently, this problem is frequently tackled with single-cell RNA sequencing (scRNA-seq) technologies by analysing messenger-RNA (mRNA) molecule counts as a measure of gene expression (Haque et al.,, 2017). The measured gene expression (also called expression levels) provide the necessary information about the nature of cells at a specific point of time, also known as cell states, as well as their changes over time. However, due to the experimental limitations of the current sequencing methods each cell gets destroyed in the measurement process and can therefore be observed only once. This situation makes it difficult to infer cell state transitions and the overall sequence of cell states of a temporal biological process. To that end, pseudotime ordering is a popular approach to describe such a biological process as a sequence of cell states along a time sequence. (Trapnell et al.,, 2014).
@@ 1.2 Contributions
  • We develop a probabilistic GP modeling framework for latent (input) variable estimation using derivative information for any multi-dimensional data-generating process. Our model accounts for dimension specific information and interactions between dimensions in a multi-dimensional data scenario which are common in fields like single-cell biology.
@@
  • We showcase the application of our modeling approach on a reduced real-world scRNA data set.
@@ 2 Related work
Gaussian processes, as a class of models, underwent a wide range of extensions over the years giving rise to various forms of GP models. Specifically, three broad extensions relevant to this work are GPs with derivative information, multi-output GPs and GPs for latent variable modeling. Using derivative information for Gaussian processes was introduced in Solak et al., (2002) who replaced standard covariance functions with their derivative counterparts. This paved the way to modeling data along with its derivatives as a single GP model. Recently, derivative GPs were extended to support multiple inputs and scalable approximations (Eriksson et al.,, 2018; Padidar et al.,, 2021). In case of multi-output GPs, recent works (Moreno-Muñoz et al.,, 2018; Joukov and Kulić,, 2022) study GPs with support for multiple outputs that are of varying nature in terms of data types, however multi-output GPs with derivative information have not been studied in detail yet. Latent GPs were introduced by Lawrence, (2003, 2005) and have, so far, been predominantly used for dimensionality reduction (Titsias,, 2009; Titsias and Lawrence,, 2010). More recently, for similar applications of dimension-reduction technique, extensions on GP latent variable model for non-Gaussian likelihoods with different types of latent input structures were discussed in Lalchand et al., (2022). These works also focus on scalable approximations to latent GPs. In contrast, we focus estimating latent inputs that probabilistically explains a dependent multi-output variable.
For the modeling of scRNA sequencing data, GPs have been broadly applied in two relevant directions, specifically, for clustering (Buettner and Theis,, 2012; Buettner et al.,, 2015) and temporal modeling (Hensman et al., 2013b, ). Considering the latter, pseudotime estimation constitutes a major research direction as it is directly related to understanding the true underlying biological processes. It has been shown previously that point estimates of pseudotime are highly prone to infer false cellular ordering, thus suggesting Bayesian inference to provide uncertainty estimates alongside each estimated pseudotime (Campbell and Yau,, 2015). Further works focus on latent pseudotime estimations (Reid and Wernisch,, 2016; Campbell and Yau,, 2018) along with branching structures for trajectory inference (Ahmed et al.,, 2019) based on a GP framework. One of the main limitations in these works lies in their restricted use of gene expressions, taking into account expression profile snapshots as the only available information regarding cellular ordering. We provide evidence that including RNA velocity as derivative information holds the power to estimate latent pseudotime with increased precision compared to what previous approaches could achieve.
@@ 3 Methods
We develop DGP-LVM, a framework for derivative Gaussian process modeling with the primary goal of estimating latent variables serving as (implicit) inputs to the GP. As the general setup, we consider a pair of variables $(y,x)$ where $y$ is the output (response) variable and $x$ is the input variable (covariate), with individual observations denoted as $y_{i},x_{i}\in\mathbb{R},i=\{1,...,N\}$ where $N$ is the number of observations. In addition to $y$ itself, we incorporate the derivative outputs $y^{\prime}=\delta y/\delta x$ into the model. The components of DGP-LVM are first discussed individually, before we combine them into a single model.
@@ 3.1 Derivative Gaussian processes
A GP is a stochastic process specified by a mean function $m(x)$ (usually fixed to zero), and a covariance function $K=K(x,x^{T})$, where $x^{T}$ indicate transpose of $x$, such that a finite set of these points will follow a multivariate Gaussian distribution (Williams and Rasmussen,, 1995). Here, we consider zero-mean GPs $f(x)$ such that $f(x)\sim\mathcal{GP}(0,K)$. If the output variable $y$ is univariate, modeling the relationship of $x$ and $y$ via a (single-output) GP and independent additive noise can be written as
@@
$\left(\begin{matrix}f(x)\\ f^{\prime}(x)\end{matrix}\right)\sim\mathcal{GP}\left(0,\left(\begin{matrix}K&K^{\prime}\\ K^{\prime T}&K^{\prime\prime}\end{matrix}\right)\right),$ (3)
where $K^{\prime}$ is the first derivative of the covariance function $K=K(x,x^{T})$ with respect to $x$ and $K^{\prime T}$ is the first derivative of the covariance function with respect to $x^{T}$. $K^{\prime\prime}$ is the second order partial derivative of $K$ differentiating both with respect to $x$ and $x^{T}$. In other words, differentiation simply propagates through the covariance function (see Appendix A for mathematical details). The specific properties of such derivative GP models depend on the chosen covariance function. A common choice is the Squared Exponential (SE) covariance function with hyperparameters $\rho$ as length scale and $\alpha$ as the GP marginal standard deviation (SD). The derivative version of the SE covariance function is given by
@@
Derivative covariance functions are obtainable generally for any chosen covariance function whose second order derivative exists. In this paper, we focus on SE and Matern class covariance functions as perhaps the most common choices. We provide derivative SE, Matern 3/2 and Matern 5/2 covariance functions in Appendix A. In our real-world case study (Section 5), we specifically consider the SE and Matern 3/2 as the extreme cases in terms of smoothness among the Matern class covariance functions.
@@ 3.1.1 Customised hyperparameters
Properly including the derivative observations $y^{\prime}$ requires more than just using a basic derivative covariance function. Due to the properties of differentiation, $y^{\prime}$ can be on a fundamentally different scale than $y$ and thus, needs to be treated as such. In addition to having a different mean (i.e., different GP components $f(x)$ vs. $f^{\prime}(x)$), the error SDs for $y$ and $y^{\prime}$ will also be different. Moreover, for real data, the observations $y^{\prime}$ containing derivative information may not be the exact same as the true derivatives $\delta y/\delta x$, but only be proportional to them (see Section 1). This proportionality induces a scale difference between $y^{\prime}$ and what is canonically modeled by a basic derivative covariance function. This creates a major issue for models ignoring scale differences as we demonstrate in our simulations.
@@
We also introduce a second error SD parameter $\sigma^{\prime}$ corresponding only to the derivative outputs. We assume both the GP marginal SD and error SD parameters are related to each other, respectively, through scalar multiples due to scale difference between $y$ and $y^{\prime}$. The scale of $\rho$ is only dependent on the scale of $x$, which is constant across outputs and their derivatives, such that $\rho$ does not need to be split up into two parameters. Together, independent of specifically chosen covariance function, the DGP-LVM on $y$ and $y^{\prime}$ with independent, additive Gaussian noise is then specified as
@@ 3.2 Multidimensional outputs
Multivariate output GPs (or multi-output GPs) model multiple response variables $\{y_{1},...,y_{D}\}$ jointly over $D>1$ output dimensions (Rasmussen and Williams,, 2006). Extending our univariate notation, the individual output values are now denoted as $y_{di}$ for dimension $d$ and observation $i$, with corresponding derivative values $y^{\prime}_{di}$. Multi-output GPs are created by first setting up $D$ independent, univariate Gaussian processes $f_{d}(x)$ each with their own set of hyperparameters, that is, $(\rho_{d},\alpha_{d}$, $\alpha^{\prime}_{d}$, $\sigma_{d}$ and $\sigma^{\prime}_{d})$ for Matern covariance functions with adjusted scales. Subsequently the univariate GPs are related to one another by folding them with a ($D$-dimensional) across-dimension correlation matrix $C$ (Teh et al.,, 2005; Bonilla et al.,, 2007). That is, for each observation $i$, we obtain a vector of across-dimension correlated GP values as
@@ 3.3 Latent variable inputs
So far, we have considered the input $x$ to be known exactly. However, in practice, we often only have a noisy measurement $\tilde{x}$ of $x$ available. In this context, the true $x$ becomes a latent variable, which needs to be appropriately modeled and subsequently estimated. If we assume that the measurements $\tilde{x}$ are Gaussian with known measurement SD $s$, we can write for each observation $i$:
@@ 3.4 The full model
Combining all of the above model components results in our DGP-LVM framework. As such, DGP-LVM can be applied to the aforementioned problem of pseudotime estimation from single-cell RNA sequencing data. The scRNA sequencing data we consider consists of spliced RNA gene counts and RNA velocity, the time derivative of gene counts. DGP-LVM allows inclusion of both these information into a single model (see Section 3.1). Since the RNA velocity is not an exact derivative of spliced RNA counts, it induces a scale difference that is solved by DGP-LVM as shown in Section 3.1.1. Given that single-cell RNA sequencing data is multi-dimensional, DGP-LVM is designed as a multi-output model (see Section 3.2). The primary aim of DGP-LVM is to estimate the latent inputs as explained in Section 3.3, which perfectly aligns with pseudotime estimation since pseudotime is an unobserved cell ordering, hence is considered as a latent variable. Moreover, the RNA sequencing data comes with its own cell capture time or experimental time which can be considered as a noisy version of the true pseudotime.
3.5 Bayesian inference
We estimate DGP-LVMs in a Bayesian framework using MCMC sampling. In DGP-LVM, the parameters of interest are the latent inputs $x$ with the data being $y$ and $y^{\prime}$ and the set of covariance function hyperparameters denoted by $\theta$. Following from the specification in (10), the joint probability density factorizes as
$p(y,y^{\prime},x,\theta)=p(y\mid x,\theta)\,p(y^{\prime}\mid x,\theta)\,p(x)\,p(\theta).$ (14)
where $p(y\mid x,\theta)$ and $p(y^{\prime}\mid x,\theta)$ denote the respective GP-based likelihoods, $p(x)$ denotes the prior for the latent $x$ implied by the measurement model (13), and $p(\theta)$ denotes priors for the covariance function hyperparameters. More details on the choice of prior distributions are discussed in Section 4.2. Using Bayes’ rule, we obtain the joint posterior over $x$ and $\theta$ as
@@ 3.5 Bayesian inference
Posterior samples of $x$ and $\theta$ are obtained through MCMC sampling via extensions of Hamiltonian Monte Carlo (Neal,, 2011; Hoffman and Gelman,, 2014). We implemented all models in Stan using the RStan interface (Stan Development Team,, 2024).

### v3

@@ Abstract
We develop a framework for derivative Gaussian process latent variable models (DGP-LVMs) that can handle multi-dimensional output data using modified derivative covariance functions. The modifications account for complexities in the underlying data generating process such as scaled derivatives, varying information across multiple output dimensions as well as interactions between outputs. Further, our framework provides uncertainty estimates for each latent variable samples using Bayesian inference. Through extensive simulations, we demonstrate that latent variable estimation accuracy can be drastically increased by including derivative information due to our proposed covariance function modifications. The developments are motivated by a concrete biological research problem involving the estimation of the unobserved cellular ordering from single-cell RNA (scRNA) sequencing data for gene expression and its corresponding derivative information known as RNA velocity. Since the RNA velocity is only an estimate of the exact derivative information, the derivative covariance functions need to account for potential scale differences. In a real-world case study, we illustrate the application of DGP-LVMs to such scRNA sequencing data. While motivated by this biological problem, our framework is generally applicable to all kinds of latent variable estimation problems involving derivative information irrespective of the field of study.
Keywords: Gaussian processes, single-cell RNA, Bayesian inference, derivative GPs, latent variables
@@ 1 Introduction
Gaussian processes (GPs) are a class of statistical models known for their flexible structure and favourable properties to analyse complex data (Williams and Rasmussen,, 1995). Since their inception, several extensions have been proposed, among which the most relevant ones to this paper are adding derivative information to GPs (Solak et al.,, 2002; Rasmussen and Williams,, 2006), building GPs with multiple outputs (Cressie,, 1993; Teh et al.,, 2005; Rasmussen and Williams,, 2006) as well as modelling latent input variables (Lawrence,, 2003, 2005). Since differentiation is a linear operation, any variable and its derivative would be linearly related. For the same reason, as a fundamental property of GPs, a derivative of a GP is just another GP with a related covariance function. Together, this results in a single GP model for the outputs and their derivatives with a joint derivative covariance function (Solak et al.,, 2002).
@@
In real-world data, the derivatives are seldom exactly computed, which adds a major challenge to the modelling endeavour. If derivatives are computed with respect to the observed (non-latent) inputs, they are naturally only an approximation of the derivatives with respect to latent inputs. Conversely, if the derivatives are (implicitly) computed with respect to the latent inputs, the uncertainty of the latter will induce hard-to-quantify uncertainty in the estimated derivatives. One way or another, this lack of exact derivative information poses a serious challenge for derivative GP modelling. Existing approaches are not equipped to deal with the significant scale differences between outputs and derivatives, thus requiring modifications in its covariance functions to ensure valid and efficient latent variable estimation.
In this paper, we demonstrate a combination of all the above model extensions leading to our DGP-LVM: derivative Gaussian process latent variable model framework. We provide more context about the real-world modelling challenges in Section 1.1 as a basis for our motivation to develop DGP-LVM. The remainder of the paper is structured as follows. We discuss and provide context on related works in Section 2. We introduce our methodology and model development in Section 3 and perform extensive simulation studies in Section 4 that demonstrate the relevance of our contributions. We further illustrate DGP-LVM on a real single-cell RNA sequencing data in Section 5 before discussing our methods’ limitations and future work in Section 6.
@@ 1.1 Motivation
In developmental biology, to describe temporal biological processes, researchers use stochastic approaches to understand cellular progression, that is, how cells develop and undergo changes in their state throughout various stages over the period of time (Maamar et al.,, 2007; Losick and Desplan,, 2008; Raj and Oudenaarden,, 2008). Currently, this problem is frequently tackled with single-cell RNA sequencing (scRNA-seq) technologies by analysing messenger-RNA (mRNA) molecule counts as a measure of gene expression (Haque et al.,, 2017). The measured gene expression (also called expression levels) provide the necessary information about the nature of cells at a specific point of time, also known as cell states, as well as their changes over time. However, due to the experimental limitations of the current sequencing methods each cell gets destroyed in the measurement process and can therefore be observed only once. This situation makes it difficult to infer cell state transitions and the overall sequence of cell states of a temporal biological process. To that end, pseudotime ordering is a popular approach to describe such a biological process as a sequence of cell states along a time sequence (Trapnell et al.,, 2014).
@@ 1.2 Contributions
  • We develop a probabilistic GP modelling framework for latent (input) variable estimation using derivative information for any multi-dimensional data-generating process. Our model accounts for dimension specific information and interactions between dimensions in a multi-dimensional data scenario which are common in (but not limited to) fields like single-cell biology.
@@
  • We showcase the application of our modelling approach on a reduced real-world scRNA-seq data set.
@@ 2 Related work
Gaussian processes, as a class of models, underwent a wide range of extensions over the years giving rise to various forms of GP models. Specifically, three broad extensions relevant to this work are GPs with derivative information, multi-output GPs and GPs for latent variable modelling. Using derivative information for Gaussian processes was introduced in Solak et al., (2002) who replaced standard covariance functions with their derivative counterparts. This paved the way to modelling data along with its derivatives as a single GP model. Recently, derivative GPs were extended to support multiple inputs and scalable approximations (Eriksson et al.,, 2018; Padidar et al.,, 2021). In case of multi-output GPs, recent works (Moreno-Muñoz et al.,, 2018; Joukov and Kulić,, 2022) study GPs with support for multiple outputs that are of varying nature in terms of data types, however multi-output GPs with derivative information have not been studied in detail yet. Latent GPs were introduced by Lawrence, (2003, 2005) and have, so far, been predominantly used for dimensionality reduction (Titsias,, 2009; Titsias and Lawrence,, 2010). More recently, for similar applications of dimension-reduction technique, extensions on GP latent variable model for non-Gaussian likelihoods with different types of latent input structures were discussed in Lalchand et al., (2022). These works also focus on scalable approximations to latent GPs. In contrast, we focus on estimating latent inputs that probabilistically explains a dependent multi-output variable.
For the modelling of scRNA-seq data, GPs have been broadly applied in two relevant directions, specifically, for clustering (Buettner and Theis,, 2012; Buettner et al.,, 2015) and temporal modelling (Hensman et al., 2013b, ). Considering the latter, pseudotime estimation constitutes a major research direction as it is directly related to understanding the true underlying biological processes. It has been shown previously that point estimates of pseudotime are highly prone to infer false cellular ordering, thus suggesting Bayesian inference to provide uncertainty estimates alongside each estimated pseudotime (Campbell and Yau,, 2015). Further works focus on latent pseudotime estimations (Reid and Wernisch,, 2016; Campbell and Yau,, 2018) along with branching structures for trajectory inference (Ahmed et al.,, 2019) based on a GP framework. One of the main limitations in these works lie in their restricted use of gene expressions, taking into account expression profile snapshots as the only available information regarding cellular ordering. We provide evidence that including RNA velocity as derivative information holds the power to estimate latent pseudotime with increased precision compared to what previous approaches could achieve.
@@ 3 Methods
We develop DGP-LVM, a framework for derivative Gaussian process modelling with the primary goal of estimating latent variables serving as (implicit) inputs to the GP. As the general setup, we consider a pair of variables $(y,x)$ where $y$ is the output (response) variable and $x$ is the input variable (covariate), with individual observations denoted as $y_{i},x_{i}\in\mathbb{R},i=\{1,...,N\}$ where $N$ is the number of observations. In addition to $y$ itself, we incorporate the derivative outputs $y^{\prime}=\delta y/\delta x$ into the model. The components of DGP-LVM are first discussed individually, before we combine them into a single model.
@@ 3.1 Derivative Gaussian processes
A GP is a stochastic process specified by a mean function $m=m(x)$, and a covariance function $K=K(x,x^{T})$, where $x^{T}$ indicate transpose of $x$, such that a finite set of these points will follow a multivariate Gaussian distribution (Williams and Rasmussen,, 1995). Concretely, we consider GPs $f(x)$ such that $f(x)\sim\mathcal{GP}(m,K)$. Here we consider a constant mean function (similar to an intercept in regression models). If the output variable $y$ is univariate, modelling the relationship of $x$ and $y$ via a (single-output) GP and independent additive noise can be written as
@@
$\left(\begin{matrix}f(x)\\ f^{\prime}(x)\end{matrix}\right)\sim\mathcal{GP}\left(\left(\begin{matrix}m_{f}\\ m_{f^{\prime}}\end{matrix}\right),\left(\begin{matrix}K&K^{\prime}\\ K^{\prime T}&K^{\prime\prime}\end{matrix}\right)\right),$ (3)
where $m_{f}$ and $m_{f^{\prime}}$ are constant mean functions corresponding to GP $f$ and it’s derivative $f^{\prime}$ respectively. $K^{\prime}$ is the first derivative of the covariance function $K=K(x,x^{T})$ with respect to $x$ and $K^{\prime T}$ is the first derivative of the covariance function with respect to $x^{T}$. $K^{\prime\prime}$ is the second order partial derivative of $K$ differentiating both with respect to $x$ and $x^{T}$. In other words, differentiation simply propagates through the covariance function (see Appendix A for mathematical details). The specific properties of such derivative GP models depend on the chosen covariance function. A common choice is the Squared Exponential (SE) covariance function with hyperparameters $\rho$ as length scale and $\alpha$ as the GP marginal standard deviation (SD). The derivative version of the SE covariance function is given by
@@
Derivative covariance functions are obtainable generally for any chosen covariance function whose second order derivative exists. In this paper, we focus on SE and Matern class covariance functions as perhaps the most common choices. We provide further details on the mathematical forms of derivative SE, Matern 3/2 and Matern 5/2 covariance functions in Appendix A.
@@ 3.1.1 Customised hyperparameters
Properly including the derivative observations $y^{\prime}$ requires more than just using a basic derivative covariance function. Due to the properties of differentiation, $y^{\prime}$ can be on a fundamentally different scale than $y$ and thus needs to be treated as such. In addition to having different signals (i.e., different GP components $f(x)$ vs. $f^{\prime}(x)$), the error SDs for $y$ and $y^{\prime}$ will also be different. Moreover, for real data, the observations $y^{\prime}$ containing derivative information may not be the exact same as the true derivatives $\delta y/\delta x$, but only be proportional to them (see Section 1). This proportionality induces a scale difference between $y^{\prime}$ and what is canonically modelled by a basic derivative covariance function. This creates a major issue for models ignoring scale differences as we demonstrate in our simulations.
@@
In other words, since $K(x_{i},x_{j})=Cov(y_{i},y_{j})$ we account for the GP marginal variance through $\alpha^{2}$ since we are only concerned with the original part of the GP $f$. In the case of $Cov(y_{i},y^{\prime}_{j})$ and $Cov(y^{\prime}_{i},y_{j})$, we compute $K^{\prime}(x_{i},x_{j})$ and thus we account for the GP marginal variance through $\alpha\alpha^{\prime}$ where $\alpha$ belongs to $f$ and $\alpha^{\prime}$ belongs to $f^{\prime}$. Finally, in the case of $Cov(y^{\prime}_{i},y^{\prime}_{j})$, we compute $K^{\prime\prime}(x_{i},x_{j})$ and are thus accounting for the GP marginal variance through $\alpha^{\prime 2}$ since we are only dealing with the derivative part of the GP $f^{\prime}$. Since, in addition to the product term $\alpha\alpha^{\prime}$ in $K^{\prime}$, we also estimate $\alpha^{2}$ and $\alpha^{\prime 2}$ separately through $K$ and $K^{\prime\prime}$, respectively, estimating both $\alpha$ and $\alpha^{\prime}$ does not cause identifiability issues.
Similarly, we define two residual standard deviation parameters $\sigma$ and $\sigma^{\prime}$, accounting for measurement noise in $y$ and $y^{\prime}$, respectively. The scale of $\rho$ is only dependent on the scale of $x$, which is constant across outputs and their derivatives, such that $\rho$ does not need to be split up into two parameters. Together, independent of specifically chosen covariance function, the DGP-LVM on $y$ and $y^{\prime}$ with independent, additive Gaussian noise is then specified as
@@ 3.2 Multidimensional outputs
Multivariate output GPs (or multi-output GPs) model multiple response variables $\{y_{1},...,y_{D}\}$ jointly over $D>1$ output dimensions (Rasmussen and Williams,, 2006). Extending our univariate notation, the individual output values are now denoted as $y_{di}$ for dimension $d$ and observation $i$, with corresponding derivative values $y^{\prime}_{di}$. Multi-output GPs are created by first setting up $D$ independent, univariate Gaussian processes $f_{d}(x)$ each with their own set of hyperparameters, that is, $(\rho_{d},\alpha_{d}$, $\alpha^{\prime}_{d}$, $\sigma_{d}$ and $\sigma^{\prime}_{d})$ for Matern class of covariance functions with adjusted scales. Subsequently the univariate GPs are related to one another by folding them with a ($D$-dimensional) across-dimension correlation matrix $C$ (Teh et al.,, 2005; Bonilla et al.,, 2007). That is, for each observation $i$, we obtain a vector of across-dimension correlated GP values as
@@ 3.3 Latent variable inputs
So far, we have considered the input $x$ to be known exactly. However, in practice, we often only have a noisy measurement $\tilde{x}$ of $x$ available. In this context, the true $x$ becomes a latent variable, which needs to be appropriately modelled and subsequently estimated. If we assume that the measurements $\tilde{x}$ are Gaussian with known measurement SD $s$, we can write for each observation $i$:
@@ 3.4 The full model
Below, we summarize all the extensions that together make up our proposed DGP-LVM framework. To shorten the notation, let us denote the vector of GP hyperparameters for dimension $d$ as $\theta_{d}$. For our considered covariance functions, $\theta_{d}$ includes the length scale $\rho_{d}$, GP marginal SDs $\alpha_{d}$ and $\alpha^{\prime}_{d}$ as well as the error SDs $\sigma_{d}$ and $\sigma^{\prime}_{D}$. Considering the SE covariance function as an example, full DGP-LVMs are then specified as follows:
$\left(\begin{matrix}f_{d}(x)\\ f^{\prime}_{d}(x)\end{matrix}\right)$ $\sim\mathcal{GP}\left(\left(\begin{matrix}m_{f}\\ m_{f^{\prime}}\end{matrix}\right),\left(\begin{matrix}K_{d}&K_{d}^{\prime}\\ K_{d}^{\prime T}&K_{d}^{\prime\prime}\end{matrix}\right)\right),$ (14)
$K_{d}(x_{i},x_{j})$ $=\alpha_{d}^{2}\exp\left(-\frac{(x_{i}-x_{j})^{2}}{2\rho_{d}^{2}}\right),$
$K^{\prime}_{d}(x_{i},x_{j})$ $=\alpha_{d}\alpha_{d}^{\prime}\frac{(x_{i}-x_{j})}{\rho_{d}^{2}}\exp\left(-\frac{(x_{i}-x_{j})^{2}}{2\rho_{d}^{2}}\right),$
$K^{\prime\prime}_{d}(x_{i},x_{j})$ $=\frac{\alpha_{d}^{\prime 2}}{\rho_{d}^{4}}(\rho_{d}^{2}-(x_{i}-x_{j})^{2})\exp\left(-\frac{(x_{i}-x_{j})^{2}}{2\rho_{d}^{2}}\right),$
$y_{di}$ $\sim\mathcal{N}(f_{d}(x_{i}),\sigma_{d}^{2}),$
$y^{\prime}_{di}$ $\sim\mathcal{N}(f^{\prime}_{d}(x_{i}),\sigma_{d}^{\prime 2}),$
$\tilde{x}_{i}$ $\sim\mathcal{N}(x_{i},s^{2}),$
$\theta_{d}$ $\sim p(\theta_{d})=p(\rho_{d})\,p(\alpha_{d})\,p(\alpha^{\prime}_{d})\,p(\sigma_{d})\,p(\sigma^{\prime}_{d}).$
Following the above specifications, after marginalizing out $f$ and $f^{\prime}$, the multi-output joint probability density factorizes as
$p(y,y^{\prime},x,\theta)=\prod_{d}^{D}p(y_{d}\mid x,\theta_{d})\,p(y^{\prime}_{d}\mid x,\theta_{d})\,p(x)\,p(\theta_{d}).$ (15)
where $p(y_{d}\mid x,\theta_{d})$ and $p(y^{\prime}_{d}\mid x,\theta_{d})$ denote the respective GP-based likelihoods for a single output dimension. $p(x)$ denotes the prior for the latent $x$ implied by the measurement model Eq.(13). More details on the choice of prior distributions are discussed in Section 4.2. Using Bayes’ rule, we obtain the joint posterior over $x$ and $\theta$ as
@@
Posterior samples of $x$ and $\theta$ (i.e. all the covariance function hyperparameters) are obtained through MCMC sampling via adaptive Hamiltonian Monte Carlo (Neal,, 2011; Hoffman and Gelman,, 2014). We implemented all models in Stan using the RStan interface (Stan Development Team,, 2024).
Implemented as above, DGP-LVMs can be applied to the aforementioned problem of pseudotime estimation from single-cell RNA sequencing data. The scRNA-seq data we consider consists of spliced RNA gene counts and RNA velocity, the time derivative of gene counts. DGP-LVM allows inclusion of both these information into a single model (see Section 3.1). Since the RNA velocity is not an exact derivative of spliced RNA counts, it induces a scale difference that is solved by DGP-LVM as shown in Section 3.1.1. Given that single-cell RNA sequencing data is multi-dimensional, DGP-LVM is designed as a multi-output model (see Section 3.2). The primary aim of DGP-LVM is to estimate the latent inputs as explained in Section 3.3, which perfectly aligns with pseudotime estimation since pseudotime is an unobserved cell ordering, and is hence considered as a latent variable. Moreover, the RNA sequencing data comes with its own cell capture time or experimental time which can be considered as a noisy version of the true pseudotime.